There are 16 questions each with 5 multiple choices. What is the probability of answering exactly 7 questions correctly?


1
Expert's answer
2021-02-28T16:41:05-0500

If there are 5 multiple answers, the probability of right answer is 0.2. So the probability to answer correctly on 7 particular questions is 0.27*0.816-7 = 0.27*0.89. And we need to multiply this by C167 - binomial coefficient, which means how many variants exist to choose 7 of 16 without order. So the probability is C167 * 0.27*0.89 = 0.0196..